OpenStudy (anonymous):

General Advice on taking tests in calculus class. What is the best way to present an answer like y = 1/4x + 7/4 y = 1/4x + 1 3/4 y = 0.25x + 1.75 And will one form or another, cost you points?

OpenStudy (sweetburger):

y=.25x+1.75 is very clear to me and no you can right it anyway you want to as long as the value isnt changing

OpenStudy (anonymous):

and in a written test, would you recommend putting all the terms in brackets? y =(1/4)(x) + (7/4) for example.. just to be safe.

OpenStudy (anonymous):

which I see, you wouldnt need if it were y=0.25x + 1.75

OpenStudy (anonymous):

thnx

OpenStudy (mle):

if the question asks it with decimals, you answer with decimals if calculator *is* allowed, you can answer with decimals if no calculator is allowed or there is nothing to give away which you should use go with the improper fractions with parenthesis :)

OpenStudy (mle):

costing points would depend on the person grading the test tho I would assume no points would be taken off unless you round a decimal too soon or if your handwriting is illegible :P

OpenStudy (nincompoop):

stick with fractions

OpenStudy (nincompoop):

\(\huge \frac{x}{4} + \frac{7}{4} = \frac{x+7}{4}\)

OpenStudy (nincompoop):

calculus will test your algebra skills

OpenStudy (usukidoll):

nin for once is right. My Calculus I professor told me that Calculus is the easy part. The hard part for many students is the Algebra because they tend to forget some of the laws or rules and end up in trouble.

OpenStudy (usukidoll):

use this y = 1/4x + 7/4

OpenStudy (anonymous):

\[y=\frac{ x }{ 4 }+\frac{ 7 }{ 4 }\] is the best and easy way! :)

OpenStudy (anonymous):

Thank you to all who answered... I wish I had caught all your responses before the test. but this will help for my next tests.